OpenStudy (anonymous):

find angle 3 measure of angle 1 = x measure of angle 2 = x-6

OpenStudy (anonymous):

is there anything else known about the angles?

OpenStudy (anonymous):

m∠3 = 180 -(x + x - 6 ) = 180 - 2x + 6 = 186 - 2x. Here is the steps to the problem: First, we solve for x. a=x b=x-6 c= c x+(x-6)+c=180 2x-6+c=180 +6 +6 2x+c=186 -c -c 2x =186-c (2x)/2 = (186-c)/2 x = 93-c a=93-c b=87-c c= c (93-c)+(87-c)+c=180 -93 -87 -180 3=0 3c/=0/3 c=0 a=93 b=87 c=0